3. Find the partial fraction decomposition of the function f(1) = work. ² + 4 13 - 3.12 +2r Show all of your

please show work because I'm confused. calc 2
Answer #1

f(x) = 13312 ५ 1-3142 +-3:2424 = [-1-31+2) -[-12-21-1+2) -1565-7)-1 (1-2) -1(1-DE-) ५+५ [T-1) (1-2) ५ - A + + (2-2/I-RR 1-R YEquating Co-efficients on both sides A+b+c=1 - -3A-2B-C =0 -2 ZA=4 => A=ų =2 [A=2] Substitute A=2 in 0 10 2+B+C=1) B+carl - 4